# Tensor Product Attention Is All You Need

### v6

@@ A.4 Inference-time decoding cost across mechanisms
GQA ($N_{g}=8$, 132M parameters):
$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=7168\cdot 128\cdot(2\cdot 64+2\cdot 8)\approx 132\times 10^{6}$
@@
$\text{Projection}=7168\cdot((512+64)\cdot 64+512+64)=268.4\times 10^{6}$
@@
$\text{Attention}=M\cdot[1(3136)+1(8256)]=11392M$
@@
$\text{Attention}=M\cdot[2(3136)+2(8256)]=22784M$
@@
$\text{Attention}=M\cdot[1(1600)+1(8256)]=9856M$
@@
$\text{Attention}=M\cdot[2(1600)+2(8256)]=19712M$
@@ Appendix B More on FlashTPA Decoding Algorithm
$\mathbf{A}_{Q}\!\in\!\mathbb{R}^{B\times N\times H\times R_{Q}},\;\mathbf{B}_{Q}\!\in\!\mathbb{R}^{B\times N\times R_{Q}\times D},\;\mathbf{A}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times H\times R_{K}},\;\mathbf{B}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{K}\times D},$
$\mathbf{A}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times H\times R_{V}},\;\mathbf{B}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{V}\times E}.$
We use scalings $s_{Q}{=}1/R_{Q}$, $s_{K}{=}1/R_{K}$, $s_{V}{=}1/R_{V}$, and $s_{\text{total}}{=}1/\sqrt{D}$. Let $\mathsf{mask}\!\in\!\{0,-\infty\}^{B\times N\times M}$ encode causality/padding. If RoPE pre‑rotation is used (Section 3.2), $\mathbf{B}_{K}^{\text{cache}}$ already includes positional phases; otherwise apply $\operatorname{RoPE}$ to $\mathbf{B}_{K}$ on load (no other change).
@@
3:Initialize $\mathbf{y}\leftarrow 0^{B\times H\times N\times E}$, $\mathbf{lse}\leftarrow 0^{B\times H\times N}$, $\mathbf{m}\leftarrow(-\infty)^{B\times H\times N}$
@@
5:  Load $\mathbf{B}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{K}\times D}$, $\mathbf{A}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times H\times R_{K}}$
6:  Load $\mathbf{A}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times H\times R_{V}}$, $\mathbf{B}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{V}\times E}$, $\mathsf{mask}_{\text{blk}}\!\in\!\mathbb{R}^{B\times N\times\Delta m}$
@@
9:  $\mathcal{L}_{\text{blk}}\leftarrow(s_{\text{total}}s_{Q}s_{K})\cdot\mathrm{einsum}(\text{``bnhr,bmhs,bnmrs$\to$bhnm''},~\mathbf{A}_{Q},~\mathbf{A}_{K,\text{blk}},~\mathbf{P})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times\Delta m}$
@@
13:  (4) Block value mixing:   $\mathbf{V}_{\text{blk}}\leftarrow\mathrm{einsum}(\text{``bmhu,bmue$\to$bhme''},~\mathbf{A}_{V,\text{blk}},~\mathbf{B}_{V,\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times H\times\Delta m\times E}$
14:  $\mathbf{y}_{\text{blk}}\leftarrow\mathrm{einsum}(\text{``bhnm,bhme$\to$bhne''},~\mathbf{p}_{\text{blk}},~\mathbf{V}_{\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times E}$
@@
17:  $\mathbf{lse}\leftarrow\exp(\mathbf{m}-\mathbf{m}_{\text{new}})\odot\mathbf{lse}~+~\exp(\mathbf{m}_{\text{blk}}-\mathbf{m}_{\text{new}})\odot\mathbf{s}_{\text{blk}}$;  $\mathbf{m}\leftarrow\mathbf{m}_{\text{new}}$
@@
19:return $\mathbf{O}\leftarrow s_{V}\cdot\frac{\mathbf{y}}{\mathbf{lse}[...,None]}$ permuted to $(B,N,H,E)$
@@ B.1 Triton FlashTPA Decoding Kernel
We implement the experiments using Triton [tillet2019triton]; Algorithm 3 sketches the kernel corresponding to Algorithm 2. The provided kernel outline specializes to the frequently used case $R_{K}{=}R_{V}{=}1$; general ranks follow by tiling over $R_{K},R_{V}$ and replacing the rank‑1 vector–matrix products with the corresponding small GEMMs in steps $S1/S2$ and the value mixing path.
@@
1:Input Tensors: $\mathbf{A}_{Q}(B,N,H,R_{Q})$, $\mathbf{a}^{K}(B,M,H)$, $\mathbf{a}^{V}(B,M,H)$, $\mathbf{B}_{Q}(B,N,R_{Q},D)$, $\bm{b}^{K}(B,M,D)$, $\bm{b}^{V}(B,M,E)$
@@
15:Load $\mathbf{A}_{Q,\text{local}}^{(R_{Q}\times B_{H})}$ from $\mathbf{A}_{Q}[b,0,h_{\text{start}}\dots,:]$
@@ B.2 Additional Experimental Results
  • MQA (green line) and GQA (red line) offer significant speedups over MHA and perform very similarly to each other, often overlapping in the plots.
  • MLA (purple line) demonstrates strong performance, generally being faster than MQA/GQA, particularly at longer sequence lengths.
  • FlashTPA shows excellent scalability. While at very short sequence lengths (e.g., $2^{12}$ to $2^{13}$), its performance is comparable to MQA/GQA and MLA, its decoding time increases at a notably slower rate with sequence length. Consequently, FlashTPA becomes significantly faster than MQA/GQA for sequences longer than approximately $2^{14}$.
@@
  • FlashTPA (blue line) remains one of the most efficient decoding methods. MHA (orange line) is consistently the slowest, while MQA (green line) and GQA (red line) offer considerable improvements over MHA.
  • MLA (purple line) and FlashTPA are the top two performers. FlashTPA consistently matches or exceeds the speed of MLA, particularly at longer sequence lengths (e.g., beyond $2^{15}$ or $2^{16}$ depending on the batch size). Its advantage often becomes more pronounced at the longest sequences tested ($2^{19}$). For instance, in batch size 1, TPA is clearly faster than MLA for sequence lengths $2^{16}$ and above. A similar trend is seen across other batch sizes, where TPA maintains a competitive edge or becomes superior at longer contexts.
@@
  • FlashTPA (blue line) is highly competitive. MHA (orange line) remains the least performant. MQA (green line) and GQA (red line) are faster than MHA.
  • At very short sequence lengths (around $2^{12}$ to $2^{13}$), MQA/GQA can be slightly faster than or comparable to TPA and MLA, especially for smaller batch sizes (e.g., Batch Size 1).
  • However, as sequence length increases, both MLA (purple line) and FlashTPA demonstrate superior scalability. FlashTPA generally matches or outperforms MLA, particularly for sequences longer than $2^{15}$. For example, with a batch size of 16, TPA shows a clear speed advantage over MLA for sequence lengths $2^{16}$ and greater.
@@ Appendix D Proofs of Theorems
$\operatorname{RoPE}(\mathbf{Q}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}).$
@@
$\operatorname{RoPE}(\mathbf{K}_{s})=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{s})^{\top}\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{s}).$
@@
$\operatorname{RoPE}_{t-s}(\mathbf{q}_{t,i})^{\top}\mathbf{k}_{s,i}=\widetilde{\mathbf{q}}_{t,i}^{\top}\widetilde{\mathbf{k}}_{s,i},$
@@
$\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i})=\mathbf{T}_{t}\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}},\quad\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i})=\mathbf{T}_{s}\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}.$
@@ Appendix E More Related Work
Transformers and Attention. As a sequence-to-sequence architecture, Transformer [vaswani2017attention] introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies. Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns [child2019generating, DBLP:conf/iclr/Shi0LR0LJ23, DBLP:conf/iclr/HanJKMWZ24, liang2024conv, li2024tighter, liang2024beyond], kernel-based projections [choromanski2020rethinking], and linearized transformers [tsai2019transformer, katharopoulos2020transformers, schlag2021linear, zhang2023trained, sun2023retentive, DBLP:conf/iclr/ZhangBKR24]. To decrease memory usage and circumvent the limitation of memory bandwidth in training, [shazeer2019fast] proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head. To tackle the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) [ainslie2023gqa] divides queries into several groups, and each group of queries shares a single key head and value head. Recently, DeepSeek-V2 [liu2024deepseek] applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value. Concurrently, [hu2024multi] proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q. Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
@@
KV Cache Optimization. During the auto-regressive inference of Transformers, key and value (KV) tensors from previous tokens are cached to avoid recomputation, a technique first proposed by [ott2019fairseq]. This Key-Value (KV) cache, while crucial for efficiency, consumes significant memory and can introduce latency bottlenecks due to memory bandwidth limitations [adnan2024keyformer]. Consequently, various studies have explored methods to mitigate these issues. These include KV cache eviction strategies that discard less significant tokens [zhang2023h2o, xiao2023efficient, cai2024pyramidkv, adnan2024keyformer], dynamic sparse attention mechanisms focusing on selected keys and values [ribar2023sparq, tang2024quest, singhania2024loki], offloading the KV cache to CPU memory [he2024fastdecode, lee2024infinigen, sun2024shadowkv], and quantizing the KV cache [xiao2023smoothquant, liu2024kivi, hooper2024kvquant]. In contrast to these approaches, TPA focuses on reducing the intrinsic size of the KV cache by employing tensor-decomposed key and value representations.
@@
Low-Rank Factorizations. Low-rank approximations are widely used to compress model parameters and reduce computational complexity. Notable examples include LoRA [hu2021lora], which factorizes weight updates during fine-tuning, and its derivatives tailored for various training scenarios such as efficient pretraining (ReLoRA [lialin2023relora], MoRA [jiang2024mora]), long-context training (LongLoRA [chen2023longlora], SinkLoRA [zhang2024sinklora]), and continual training (InfLoRA [liang2024inflora], GS-LoRA [zhao2024continual], I-LoRA [ren2024analyzing]). These methods generally produce static low-rank expansions that are independent of the input context. Theoretical justifications for the expressiveness of low-rank approximations have been provided by [malladi2023kernel, zeng2023expressive]. Initialization strategies for these factorization matrices have also been explored: OLoRA [buyukakyuz2024olora] utilizes QR-decomposition of pretrained weights for improved language model performance, while LoLDU [shi2024loldu] employs LDU-decomposition to accelerate LoRA training. Furthermore, AdaLoRA [zhang2023adalora] uses Singular Value Decomposition (SVD) on pretrained weights and introduces parameter importance scores to dynamically adjust ranks. TPA, in contrast, constructs Q, K, and V tensors using contextually-aware factorizations, allowing for dynamic adaptation based on the input.

### v7

@@ A.4 Inference-time decoding cost across mechanisms
GQA ($G=8$, 132M parameters):
$\text{Parameters}=d_{\text{model}}d_{h}(2H+2G)=7168\cdot 128\cdot(2\cdot 64+2\cdot 8)\approx 132\times 10^{6}$
@@
$\text{Projection}\approx d_{\text{model}}d_{c}^{\prime}+d_{c}^{\prime}H(d_{h}+d_{h}^{R})+d_{\text{model}}(d_{c}+d_{h}^{R})+Hd_{c}d_{h}$
$\hskip 42.19432pt=7168\cdot 1536+1536\cdot 64\cdot(128+64)+7168\cdot(512+64)+64\cdot 512\cdot 128$
$\hskip 42.19432pt\approx 38.2\times 10^{6}$
@@
$\text{Attention}=M\cdot[1(3072)+1(8192)]=11264M$
@@
$\text{Attention}=M\cdot[2(3072)+2(8192)]=22528M$
@@
$\text{Attention}=M\cdot[1(1536)+1(8192)]=9728M$
@@
$\text{Attention}=M\cdot[2(1536)+2(8192)]=19456M$
@@ Appendix B More on FlashTPA Decoding Algorithm
$\mathbf{A}_{Q}\!\in\!\mathbb{R}^{B\times N\times R_{Q}\times H},\;\mathbf{B}_{Q}\!\in\!\mathbb{R}^{B\times N\times R_{Q}\times D},\;\mathbf{A}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{K}\times H},\;\mathbf{B}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{K}\times D},$
$\mathbf{A}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{V}\times H},\;\mathbf{B}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{V}\times E}.$
We use scalings $s_{Q}{=}1/R_{Q}$, $s_{K}{=}1/R_{K}$, $s_{V}{=}1/R_{V}$, and $s_{\text{total}}{=}1/\sqrt{D}$. Let $\mathsf{mask}\!\in\!\{0,-\infty\}^{B\times N\times M}$ encode causality/padding. If RoPE pre‑rotation is used (Section 3.2), $\mathbf{B}_{K}^{\text{cache}}$ already includes positional phases; otherwise apply $\operatorname{RoPE}$ to $\mathbf{B}_{K}$ on load.
@@
3:Initialize $\mathbf{y}\leftarrow 0^{B\times H\times N\times E}$, $\mathbf{s}\leftarrow 0^{B\times H\times N}$, $\mathbf{m}\leftarrow(-\infty)^{B\times H\times N}$ $\triangleright$ $\mathbf{s}$ accumulates $\sum\exp(\cdot)$; log-sum-exp is $\log\mathbf{s}+\mathbf{m}$
@@
5:  Load $\mathbf{B}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{K}\times D}$, $\mathbf{A}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{K}\times H}$
6:  Load $\mathbf{A}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{V}\times H}$, $\mathbf{B}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{V}\times E}$, $\mathsf{mask}_{\text{blk}}\!\in\!\mathbb{R}^{B\times N\times\Delta m}$
@@
9:  $\mathcal{L}_{\text{blk}}\leftarrow(s_{\text{total}}s_{Q}s_{K})\cdot\mathrm{einsum}(\text{``bnrh,bmsh,bnmrs$\to$bhnm''},~\mathbf{A}_{Q},~\mathbf{A}_{K,\text{blk}},~\mathbf{P})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times\Delta m}$
@@
13:  (4) Block value aggregation (fused over $m,u$):
14:  $\mathbf{y}_{\text{blk}}\leftarrow\mathrm{einsum}(\text{``bhnm,bmuh,bmue$\to$bhne''},~\mathbf{p}_{\text{blk}},~\mathbf{A}_{V,\text{blk}},~\mathbf{B}_{V,\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times E}$
@@
17:  $\mathbf{s}\leftarrow\exp(\mathbf{m}-\mathbf{m}_{\text{new}})\odot\mathbf{s}~+~\exp(\mathbf{m}_{\text{blk}}-\mathbf{m}_{\text{new}})\odot\mathbf{s}_{\text{blk}}$;  $\mathbf{m}\leftarrow\mathbf{m}_{\text{new}}$
@@
19:return $\mathbf{O}\leftarrow s_{V}\cdot\frac{\mathbf{y}}{\mathbf{s}[...,None]}$ permuted to $(B,N,H,E)$
@@ B.1 Triton FlashTPA Decoding Kernel
We implement the experiments using Triton [57]; Algorithm 3 sketches the kernel corresponding to Algorithm 2. The provided kernel outline specializes to the frequently used case $R_{K}{=}R_{V}{=}1$; general ranks follow by tiling over $R_{K},R_{V}$ and replacing the rank‑1 vector–matrix products with the corresponding small GEMMs in steps $S1/S2$ and the value mixing path.
@@
1:Input Tensors: $\mathbf{A}_{Q}(B,N,R_{Q},H)$, $\mathbf{a}^{K}(B,M,H)$, $\mathbf{a}^{V}(B,M,H)$, $\mathbf{B}_{Q}(B,N,R_{Q},D)$, $\bm{b}^{K}(B,M,D)$, $\bm{b}^{V}(B,M,E)$
@@
15:Load $\mathbf{A}_{Q,\text{local}}^{(R_{Q}\times B_{H})}$ from $\mathbf{A}_{Q}[b,0,:,h_{\text{start}}\dots]$
@@ B.2 Additional Experimental Results
  • MQA (purple line) and GQA (green line) offer significant speedups over MHA and perform very similarly to each other, often overlapping in the plots.
  • MLA (blue line) demonstrates strong performance, generally being faster than GQA, particularly at longer sequence lengths.
  • FlashTPA shows excellent scalability. While at very short sequence lengths (e.g., $2^{12}$ to $2^{13}$), its performance is comparable to MQA/GQA and MLA, its decoding time increases at a notably slower rate with sequence length. Consequently, FlashTPA becomes significantly faster than GQA for sequences longer than approximately $2^{14}$.
@@
  • FlashTPA (red line) remains one of the most efficient decoding methods. MHA (orange line) is consistently the slowest, while MQA (purple line) and GQA (green line) offer considerable improvements over MHA.
  • MLA (blue line) and FlashTPA are the top two performers. FlashTPA consistently matches or exceeds the speed of MLA, particularly at longer sequence lengths (e.g., beyond $2^{15}$ or $2^{16}$ depending on the batch size). Its advantage often becomes more pronounced at the longest sequences tested ($2^{19}$). For instance, in batch size 1, TPA is clearly faster than MLA for sequence lengths $2^{16}$ and above. A similar trend is seen across other batch sizes, where TPA maintains a competitive edge or becomes superior at longer contexts.
@@
  • FlashTPA (red line) is highly competitive. MHA (orange line) remains the least performant. MQA (purple line) and GQA (green line) are faster than MHA.
  • However, as sequence length increases, both MLA (blue line) and FlashTPA demonstrate superior scalability. FlashTPA generally matches or outperforms MLA, particularly for sequences longer than $2^{15}$. For example, with a batch size of 16, TPA shows a clear speed advantage over MLA for sequence lengths $2^{16}$ and greater.
@@ Appendix D Proofs of Theorems
$\operatorname{RoPE}_{t}(\mathbf{Q}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}).$
@@
$\operatorname{RoPE}_{s}(\mathbf{K}_{s})=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{s})^{\top}\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{s}).$
@@
$(\mathbf{q}_{t,i}\mathbf{T}_{t-s})\,\mathbf{k}_{s,i}^{\top}=(\mathbf{q}_{t,i}\mathbf{T}_{t})\,(\mathbf{k}_{s,i}\mathbf{T}_{s})^{\top},$
@@
$\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}_{t}(\mathbf{q}_{t,i})=\mathbf{q}_{t,i}\mathbf{T}_{t}\in\mathbb{R}^{1\times d_{h}},\quad\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}_{s}(\mathbf{k}_{s,i})=\mathbf{k}_{s,i}\mathbf{T}_{s}\in\mathbb{R}^{1\times d_{h}}.$
@@ Appendix E More Related Work
Transformers and Attention. As a sequence-to-sequence architecture, Transformer [60] introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies. Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns [10, 49, 16, 30, 27, 31], kernel-based projections [11], and linearized transformers [59, 25, 44, 69, 54, 67]. To decrease memory usage and circumvent the limitation of memory bandwidth in training, [46] proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head. To tackle the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) [2] divides queries into several groups, and each group of queries shares a single key head and value head. Recently, DeepSeek-V2 [32] applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value. Concurrently, [21] proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q. Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
@@
KV Cache Optimization. During the auto-regressive inference of Transformers, key and value (KV) tensors from previous tokens are cached to avoid recomputation, a technique first proposed by [40]. This Key-Value (KV) cache, while crucial for efficiency, consumes significant memory and can introduce latency bottlenecks due to memory bandwidth limitations [1]. Consequently, various studies have explored methods to mitigate these issues. These include KV cache eviction strategies that discard less significant tokens [70, 62, 8, 1], dynamic sparse attention mechanisms focusing on selected keys and values [42, 55, 50], offloading the KV cache to CPU memory [17, 26, 53], and quantizing the KV cache [61, 34, 19]. In contrast to these approaches, TPA focuses on reducing the intrinsic size of the KV cache by employing tensor-decomposed key and value representations.
@@
Low-Rank Factorizations. Low-rank approximations are widely used to compress model parameters and reduce computational complexity. Notable examples include LoRA [20], which factorizes weight updates during fine-tuning, and its derivatives tailored for various training scenarios such as efficient pretraining (ReLoRA [28], MoRA [22]), long-context training (LongLoRA [9], SinkLoRA [66]), and continual training (InfLoRA [29], GS-LoRA [71], I-LoRA [41]). These methods generally produce static low-rank expansions that are independent of the input context. Theoretical justifications for the expressiveness of low-rank approximations have been provided by [38, 65]. Initialization strategies for these factorization matrices have also been explored: OLoRA [7] utilizes QR-decomposition of pretrained weights for improved language model performance, while LoLDU [48] employs LDU-decomposition to accelerate LoRA training. Furthermore, AdaLoRA [68] uses Singular Value Decomposition (SVD) on pretrained weights and introduces parameter importance scores to dynamically adjust ranks. TPA, in contrast, constructs Q, K, and V tensors using contextually-aware factorizations, allowing for dynamic adaptation based on the input.